OpenStudy (anonymous):

Maria's family drove 140 miles to her grandparents' house and averaged 56 mi/h on the way there. On the return trip, they averaged 50 mi/ hr. What was the total time Maria's family spent driving to and from her grandparents' house? A. 2.5 h B. 2.6 h C. 5.2 h D. 5.3 h

OpenStudy (anonymous):

@DoritosHero @ayyookyndall @aaronq

OpenStudy (anonymous):

Time = Distance/Speed So: Time to grandparents = 140/56 Time back home = 140/50 Total time = 140/56 + 140/50 = 5.3 So the answer is D.

OpenStudy (anonymous):

thank u and love the name

OpenStudy (anonymous):

np and thanks :>